OpenStudy (anonymous):

Find the x- and y- intercepts of 3x - 4y = 12 Help on how to do

OpenStudy (anonymous):

divide by 12 \[\frac{ x }{ 4 }+\frac{ y }{ -3 }=1\] x and y intercepts are 4,-3 in \[\frac{ x }{ a }+\frac{ y }{ b }=1\] x and y intercepts are a,b

OpenStudy (anonymous):

Can you elaborate?

OpenStudy (anonymous):

if x=0,y=b if y=0,x=a